# Entanglement entropy bounds for pure states of rapid decorrelation

### v1

@@ 2 State approximations of reduced complexity
$\mathrm{(II)}\qquad$ $\theta(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})=\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})\ +\ \gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})\,,$
for some $\alpha:\Omega_{A}\times\Omega_{B}\to\mathbb{R}$ and $\gamma:\Omega_{C}\times\Omega_{B}\to\mathbb{R}$. That is, conditioned on $\boldsymbol{\sigma}_{B}$, the configurations $\boldsymbol{\sigma}_{A}$ and $\boldsymbol{\sigma}_{C}$ are independently distributed (I), and the contributions of $\boldsymbol{\sigma}_{A}$ and $\boldsymbol{\sigma}_{C}$ to the phase function are additive (II). If both conditions hold, then
$|\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\rangle=e^{i\gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})}\Big{[}\sum_{\boldsymbol{\sigma}_{A}}e^{i\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})}\sqrt{p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B})}\,|\boldsymbol{\sigma}_{A}\rangle\Big{]}\,=:e^{i\gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})}|\phi_{A}(\boldsymbol{\sigma}_{B})\rangle\,,$ (2.5)
and thus (2.2) reduces to the lower-rank expression
$\varrho_{A}(\psi)=\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ |\phi_{A}(\boldsymbol{\sigma}_{B})\rangle\langle\phi_{A}(\boldsymbol{\sigma}_{B})|\,.$ (2.6)
The above observation leads us to consider as potential approximations of $|\psi\rangle$ state vectors of the form
$\boxed{|\psi_{(B)}\rangle:=\sum_{\boldsymbol{\sigma}}e^{i[\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})\ +\ \gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})]}\sqrt{p(\boldsymbol{\sigma}_{B})\,p(\boldsymbol{\sigma}_{C}\,|\,\boldsymbol{\sigma}_{B})\,p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B})}\ |\boldsymbol{\sigma}\rangle\,,}$ (2.7)
with phase functions $\alpha:\Omega_{A}\times\Omega_{B}\to\mathbb{R}$, $\gamma:\Omega_{C}\times\Omega_{B}\to\mathbb{R}$, which will be variationally selected, and the conditional probabilities of the given state $|\psi\rangle$. One should note the absence of $\boldsymbol{\sigma}_{C}$ in the last conditional probability term in (2.7), in which it differs from that of $|\psi\rangle$. Thus the approximation generically involves changes in both the phase and the amplitude.
@@
For the modified vectors the assumptions (I) and (II) are satisfied, and consequently $\varrho_{A}(\psi_{(B)})$ is given by (2.6). Hence, in contrast to $\varrho_{A}(\psi)$, the induced state $\varrho_{A}(\psi_{(B)})$ is guaranteed to satisfy
$S\big{(}\varrho_{A}(\psi_{(B)})\big{)}\leq\ln\big{(}\nu^{|B|}\big{)}\propto l\,|\ \partial A|\,,$ (2.8)
since the rank ($\mathrm{rank}\varrho_{A}(\psi_{(B)})$) is bounded by the number of terms in the sum in (2.6).
Remark 2.1.
@@
For concreteness sake, it may be noted that a simple way to split an additive phase as a sum of the above form is to pick somewhat arbitrarily a configuration $\boldsymbol{\eta}\in\Omega_{W}$ and set $\alpha$ and $\gamma$ as
$\begin{split}\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})&=\theta(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\eta}_{C})-\theta(\boldsymbol{\eta}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\eta}_{C})]\\ \gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})&=\theta(\boldsymbol{\eta}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\,.\end{split}$ (2.9)
@@
3 Fidelity bounds
The next question to consider is the degree to which $\varrho_{A}(\psi_{(B)})$ approximates the state of our original interest $\varrho_{A}(\psi\big{)}$. We recall that states’ proximity in trace-norm (denoted $\|\cdot\|_{1}$) is related, via the Fuchs-van de Graaf inequalities [31], to their fidelity $F(\varrho,\sigma):=\left\|\sqrt{\varrho}\sqrt{\sigma}\right\|_{1}^{2}$:
$1-\sqrt{F(\varrho,\sigma)}\leq\tfrac{1}{2}\|\varrho-\sigma\|_{1}\leq\sqrt{1-F(\varrho,\sigma)}\,$ (3.1)
In case the two states in $A$ are restrictions of a pair of pure states, by Uhlmann’s principle [52] the fidelity can be estimated through the progenitor state’s global overlap:
$1-F(\varrho_{A}(\psi),\varrho_{A}(\psi_{(B)})\leq 1-\left|\langle\psi|\psi_{(B)}\rangle\right|^{2}\,.$ (3.2)
Our upper bounds on the right side will involve two terms, quantifying the deviation of $|\psi\rangle$ from the two idealized conditions listed above as (I) and (II), which do hold for $|\psi_{(B)}\rangle$.
(I)  To address the degree of the correlations in $p(\boldsymbol{\sigma})$ between $\boldsymbol{\sigma}_{A}$ and $\boldsymbol{\sigma}_{C}$, when both are conditioned on the configuration $\boldsymbol{\sigma}_{B}$ of the buffer, we shall use the following terminology and notation.
@@ 3 Fidelity bounds
Let us note that for states $|\psi\rangle$, whose phase is an additive function of the form (1.9), the term $\vartheta_{B}(A|C)$ vanishes if the buffer’s width exceeds the range of the kernel $K$. More generally, $\vartheta_{B}(A|C)$ should be small in case the phase $\theta(\boldsymbol{\sigma})$ is presentable as a sum of quasi local terms similar to (1.9), once the buffer’s width exceeds the relevant coherence length by a factor of order $\ln|\partial A|$.
@@
Our general result will be derived through repeated applications of the following fidelity bound.
@@
$1-|\langle\psi|\psi_{(B)}\rangle|^{2}\leq 2\ {\delta}_{B}(A\,|\,C)+2\ \vartheta_{B}(A|C)\,.$ (3.7)
@@
We start with the estimate
$\frac{1}{2}\left(1-|\langle\psi|\psi_{(B)}\rangle|^{2}\right)\leq\left|1-\langle\psi|\psi_{(B)}\rangle\right|\,.$
Expressing $\psi(\boldsymbol{\sigma})=\sqrt{p(\boldsymbol{\sigma})}\exp\left(i\theta(\boldsymbol{\sigma})\right)$ and $p(\boldsymbol{\sigma})=p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{C}\boldsymbol{\sigma}_{B})\,\,p(\boldsymbol{\sigma}_{C}|\boldsymbol{\sigma}_{B})\,\,p(\boldsymbol{\sigma}_{B})$ in terms of conditional probabilities, we have
@@
$\leq\Big{(}\sum_{\boldsymbol{\sigma}}p(\boldsymbol{\sigma})\left|U_{\alpha,\gamma}(\boldsymbol{\sigma},\boldsymbol{\eta})-1\right|^{2}\Big{)}^{1/2}\Big{(}\sum_{\boldsymbol{\sigma}}p(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B})\Big{)}^{1/2}\ =\ \vartheta_{B}(A|C)\,.$ (3.10)
This yields the second term in the right side of (3.7). ∎
@@ 4 Entropic implications of rapid decorrelation
In this section, we establish two main results: 1. an area-law bound on the entanglement entropy (1.1) of a state $|\psi\rangle$ restricted to $A\subset W$, and 2. a closely related upper bound for the entanglement entropy difference of $|\psi\rangle$ and $|\psi_{(B)}\rangle$.
4.1 Bounding the density operator’s small eigenvalues
Inspecting (1.1) one may note that the entropy of any state $\varrho_{A}$ is strongly affected by the eigenvalues at the lower end of its spectrum. This is conveniently quantified by the distribution-type function
@@ 4.1 Bounding the density operator’s small eigenvalues
where the sum is over the eigenvalues of $\varrho_{A}$ labeled in decreasing order. Of particular relevance is the rate at which $\tau_{\varrho_{A}}(N)$ vanishes as $N$ increases towards the dimension of the relevant Hilbert space. The following main lemma estimates this quantity through a combination of (3.2) and the fidelity estimate (3.7).
@@
$\tau_{\varrho_{A}}(N)=1-\max\left\{\mathrm{tr}\varrho_{A}P_{N}\ |\ P_{N}\,\mbox{orthogonal projection of $\mathrm{rank}\ P_{N}\leq N$}\right\}$
$=1-\max\left\{F(\rho_{A},\widehat{\varrho}_{A})\ |\ \mbox{$\widehat{\varrho}_{A}$ is a state on $A$ with $\mathrm{rank}\ \widehat{\varrho}_{A}\leq N$}\right\}.$ (4.3)
The last equality is based on the following bound of the fidelity of the state $\varrho_{A}$ with a state $\widehat{\varrho}_{A}$ of $\mathrm{rank}\ \widehat{\varrho}_{A}\leq N$. Denoting by $P_{N}$ the orthogonal projection onto the range of $\widehat{\varrho}_{A}$, the Cauchy-Schwarz inequality for the trace norm implies
$\left\|\sqrt{\rho_{A}}\sqrt{\widehat{\varrho}_{A}}\right\|_{1}^{2}=\left\|\sqrt{\rho_{A}}P_{N}\sqrt{\widehat{\varrho}_{A}}\right\|_{1}^{2}\leq\mathrm{tr}\rho_{A}P_{N}\,\mathrm{tr}\widehat{\varrho}_{A}=\mathrm{tr}\rho_{A}P_{N}\,.$
Clearly, the inequality is saturated in case $\widehat{\varrho}_{A}$ is proportional to $\varrho_{A}$ times the projection onto its $N$ highest eigenvalues. This establishes (4.1). The stated inequality follows by combining it with the fidelity bound of Theorem 3.1. ∎
The bound (4.2) gives relevance to the following terminology. In what follows, we restrict attention to regular $A\subset W$ in the sense that the volume of a buffer $B_{l}=\{u\in W\backslash A\ |\ d(u,A)\leq l\}$ of arbitrary width $l\in\mathbb{N}$ is bounded by the surface area of $A$ times $l$, i.e.
$c_{d}\ l\ |\partial A|\leq\left|B_{l}\right|\leq C_{d}\ l\ |\partial A|$ (4.4)
with some dimension dependent $c_{d},C_{d}\in(0,\infty)$.
@@
The distance function $l_{0}$ plays the role of the length scale at which the $\varphi$-decay sets in. This could be independent of $A$, in which case $l_{0}$ is constant.
4.2 Area-law bound
A single scale application of (4.2) does not yet produce the desired entropy bound since, due to the presence of the logarithm in (1.1), the von Neumann entropy $S(\varrho_{A})$ is not a trace-norm continuous function of the state. A degree of its continuity is provided by the following variant of Fannes’ continuity bound [28, 11], which involves also rank considerations
For the reader’s convenience, a proof is included here in Appendix A.
Taking a cure from [1, 16, 8], we work with a sequence of approximating states using buffers $B_{l}$ of increasing width $l\in\mathbb{N}$. Increase in the buffer size allows to prod lower reaches of the spectrum of $\varrho_{A}(\psi)$. By such means we derive the following general entanglement bound.
@@ 4.2 Area-law bound
by the monotonicity of $F(x)=x\left(\ln(a/x)+1\right)$ for $0\leq x\leq 1\leq a$. Collecting the contributions to $S(\varrho_{A})=S_{0}+\sum_{k\geq 1}S_{k}$ and noticing that $k$ does not exceed $L(A)$, since the total number of eigenvalues of $\varrho_{A}\equiv\varrho_{A}(\psi)$ is $\nu^{|A|}$, yields the result. ∎
In order to elucidate this general statement, let us add some comments and formulate a simple consequence. In case $A$ is a cube, it is regular and $L(A)$ is proportional to the cube’s edge length. Moreover, power law decay of $\varphi$ with $\alpha>2$ is enough to ensure the convergence of the series corresponding to the sums in (4.8). One may therefore upper bound these sums independently of $A$. We hence arrive at the following statement, which is one of the main results of this work.
@@
For a state $|\psi\rangle$, which is conditionally $\varphi$-decoupled with power law decay with $\alpha>2$ beyond distance $l_{0}$, there is $C\in(0,\infty)$, which is independent of $l_{0}$, such that the entanglement entropy corresponding to any non-empty, regular $A\subset W$ is bounded by
$S(\varrho_{A}(\psi))\leq C\ |\partial A|\ (1+l_{0}(A))\,.$ (4.11)
@@
An application of Theorem 4.3 yields the claim. Note that in case $A\neq\emptyset$, the additive constants can be absorbed into the multiplicative constant in the right side of (4.11). ∎
The bound (4.11) describes a strict area law in case the decoupling distance $l_{0}(A)$ is independent of $A$. This applies in particular to stoquastic states with $p$ any (even critical) classical Gibbs probability measure corresponding to a finite range interaction, cf. (1.8). Such states fall into the category of tensor-network states, for which the area-law of the entanglement entropy is well known (cf. [53, 55]). For ground-states of systems such as the quantum Ising model, existing techniques yield $l_{0}(A)\propto\ln|\partial A|$ for general regular $A$, which results in a logarithmically corrected area law unless $d=1$. In case $W$ is a cylindrical set, $W=V\times[1,...,L]$, and $A=V\times\widehat{I}$ with $\widehat{I}\subset I$, transfer matrix techniques enable one to show $l_{0}(A)\propto\xi$ and hence a strict area law.
Our results have some, but only partial, overlap with the following rigorous results that were formulated for different classes of states, under different assumptions and restrictions.
@@
    The area-law conjecture for ground states is settled for $d=1$ as well as for exactly solvable or non-interacting systems (cf. [27] for a review). In a trailblazing work, Hastings [35] gave a proof for ground states of gapped local Hamiltonians defined over $W\subset\mathbb{Z}$. As emphasized in [16], this proof relies on the decay of correlations of $|\psi\rangle$.

### v2

@@ 2.2 Markovian state approximation
$\mkern-160.0mu\mathrm{(II)}\mkern 80.0mu\widehat{\theta}(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})=\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})+\gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})$ (2.8)
with some $\alpha:\Omega_{A}\times\Omega_{B}\to\mathbb{R}$ and $\gamma:\Omega_{C}\times\Omega_{B}\to\mathbb{R}$ which may be variationally selected. With such a choice a markovian approximation takes the form
$\boxed{|\psi_{(B)}\rangle:=\sum_{\boldsymbol{\sigma}}\sqrt{p(\boldsymbol{\sigma}_{B})}\,|\phi_{A}(\boldsymbol{\sigma}_{B})\rangle\otimes|\boldsymbol{\sigma}_{B}\rangle\otimes|\phi_{C}(\boldsymbol{\sigma}_{B})\rangle}$ (2.9)
with normalized vectors
$|\phi_{A}(\boldsymbol{\sigma}_{B})\rangle$ $:=\sum_{\boldsymbol{\sigma}_{A}}e^{i\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})}\sqrt{p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B})}\,|\boldsymbol{\sigma}_{A}\rangle$ (2.10)
$|\phi_{C}(\boldsymbol{\sigma}_{B})\rangle$ $:=\sum_{\boldsymbol{\sigma}_{C}}e^{i\gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})}\sqrt{p(\boldsymbol{\sigma}_{C}\,|\,\boldsymbol{\sigma}_{B})}\,|\boldsymbol{\sigma}_{C}\rangle\,.$
For the vector (2.9) the state has the Markov property in the sense that conditioned on $\boldsymbol{\sigma}_{B}$ it factorizes into a product state over $\mathcal{H}_{A}\otimes\mathcal{H}_{C}$. Its reduced state on $A$,
$\varrho_{A}(\psi_{(B)})=\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ |\phi_{A}(\boldsymbol{\sigma}_{B})\rangle\langle\phi_{A}(\boldsymbol{\sigma}_{B})|\,,$ (2.11)
@@
has a rank bounded by $\dim\mathcal{H}_{B}$. Hence, in contrast to $\varrho_{A}(\psi)$, the induced state $\varrho_{A}(\psi_{(B)})$ is guaranteed to satisfy
$S\big{(}\varrho_{A}(\psi_{(B)})\big{)}\leq\ln\big{(}\nu^{|B|}\big{)}\propto l\,|\ \partial A|\,.$ (2.12)
@@
Remark:  For concreteness sake, it may be noted that a simple way to split an additive phase as a sum of the above form is to pick somewhat arbitrarily a configuration $\boldsymbol{\eta}\in\Omega_{W}$ and set $\alpha$ and $\gamma$ as
$\begin{split}\alpha(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B})&=\theta(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\eta}_{C})-\theta(\boldsymbol{\eta}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\eta}_{C})\\ \gamma(\boldsymbol{\sigma}_{C},\boldsymbol{\sigma}_{B})&=\theta(\boldsymbol{\eta}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\,.\end{split}$ (2.13)
@@
3 Fidelity bound for the Markovian state approximation
Quantification of the deviation of $|\psi\rangle$ from the idealized conditions listed above as (I) and (II), which hold for $|\psi_{(B)}\rangle$, will involve two terms.
(I)  To address the degree of correlations in $p(\boldsymbol{\sigma})$ between $\boldsymbol{\sigma}_{A}$ and $\boldsymbol{\sigma}_{C}$, when both are conditioned on the configuration $\boldsymbol{\sigma}_{B}$ of the buffer, we use the following terminology and notation.
@@ 3 Fidelity bound for the Markovian state approximation
Let us note that for states $|\psi\rangle$, whose phase is an additive function of the form (1.9), the term $\vartheta_{B}(A|C)$ vanishes if the buffer’s width exceeds the range of the kernel $\widehat{J}$. More generally, $\vartheta_{B}(A|C)$ should be small in case the phase $\theta(\boldsymbol{\sigma})$ is presentable as a sum of quasi local terms similar to (1.9), once the buffer’s width exceeds the relevant coherence length by a factor of order $\ln|\partial A|$.
@@
The above quantities will next be used to estimate the degree to which $\varrho_{A}(\psi_{(B)})$ approximates the state of our original interest $\varrho_{A}(\psi)$. Their trace-norm distance relates to the fidelity as measured by the progenitor state’s global overlap (somewhat in the spirit of Uhlmann’s bound [52]):
$\|\varrho_{A}(\psi)-\varrho_{A}(\psi_{(B)})\|_{1}$ $=\sup_{\|O_{A}\|\leq 1}\left|\langle\psi|O_{A}\otimes\mathbbm{1}|\psi\rangle-\langle\psi_{(B)}|O_{A}\otimes\mathbbm{1}|\psi_{(B)}\rangle\right|$
$\leq 2\left\|\psi-\psi_{(B)}\right\|\leq 2\sqrt{2(1-\operatorname{Re\,}\langle\psi|\psi_{(B)}\rangle)}\,.$ (3.5)
Our main estimate of the fidelity of the approximation of $\rho_{A}(\psi)$ by $\rho_{A}(\psi_{(B)})$ at different choices of the buffer $B$ is the following result.
@@
$\Big{[}\tfrac{1}{2}\|\varrho_{A}(\psi)-\varrho_{A}(\psi_{(B)})\|_{1}\Big{]}^{2}\leq 2\left|1-\langle\psi|\psi_{(B)}\rangle\right|\leq 2\ {\delta}_{B}(A\,|\,C)+2\ \vartheta_{B}(A|C)\,.$ (3.6)
@@
By (3) it suffices to prove the second inequality. Expressing $\psi(\boldsymbol{\sigma})=\sqrt{p(\boldsymbol{\sigma})}\exp\left(i\theta(\boldsymbol{\sigma})\right)$ and $p(\boldsymbol{\sigma})=p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{C}\boldsymbol{\sigma}_{B})\,\,p(\boldsymbol{\sigma}_{C}|\boldsymbol{\sigma}_{B})\,\,p(\boldsymbol{\sigma}_{B})$ in terms of conditional probabilities, we have
@@
$\leq\Big{(}\sum_{\boldsymbol{\sigma}}p(\boldsymbol{\sigma})\left|U_{\alpha,\gamma}(\boldsymbol{\sigma})-1\right|^{2}\Big{)}^{1/2}\Big{(}\sum_{\boldsymbol{\sigma}}p(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B})\Big{)}^{1/2}\ =\ \vartheta_{B}(A|C)\,.$ (3.9)
This yields the second term on the right side of (3.6). ∎
Remark:  Alternatively, one could have arrived at (3.6) by starting from (2.5) using the bound $p(\boldsymbol{\sigma}_{B})\ \lambda(\boldsymbol{\sigma}_{B})\geq\left|\langle\phi_{A}(\boldsymbol{\sigma}_{B})\otimes\boldsymbol{\sigma}_{B}\otimes\phi_{C}(\boldsymbol{\sigma}_{B})|\psi\rangle\right|^{2}$ with the vectors from (2.10). Applying the estimates in the proof of Theorem 3.1 to the scalar product, one then concludes the outer estimate in (3.6).
@@ 4 Entropic implications of rapid decorrelation
In this section, we relate the fidelity lower bounds formulated above with the following entropy estimates:
  1. i.
    An area-law bound corrected by $\ln|A|$ on the entanglement entropy in pure states of rapid decorrelation in dimensions $d\geq 1$,
  2. ii.
    An improved estimate in which the log correction is replaced by $\ln|\partial A|$.
  3. iii.
    A related upper bound for the difference in entanglement entropy between $|\psi\rangle$ and $|\psi_{(B)}\rangle$.
The difference between $\ln|A|$ and $\ln|\partial A|$ accomplished in ii is of particular value for one dimension, where the latter is just a constant. It also affects iii.
Throughout the discussion we restrict attention to sets $A$ which are regular in the following sense.
Definition 4.1.
We call $A\subset W$ regular if the volume of a buffer $B_{l}:=\{u\in W\backslash A\ |\ d(u,A)\leq l\}$ of width $1\leq l\leq L(A):=|A|/(c_{d}\ |\partial A|)$ is bounded by the surface area of $A$ times $l$, i.e.
$c_{d}\ l\ |\partial A|\leq\left|B_{l}\right|\leq C_{d}\ l\ |\partial A|$ (4.1)
with some dimension dependent $c_{d},C_{d}\in(0,\infty)$.
4.1 An entanglement bound based on a single buffer estimate
In the single scale version of our bound, we approximate the state $\varrho_{A}(\psi)$ by $\varrho_{A}(\psi_{(B)})$, that is, the reduced state of the conditionally decorrelated vector (2.9). The entropy of $\varrho_{A}(\psi_{(B)})$ is upper bounded by through its Schmidt norm:
$S(\varrho_{A}(\psi_{(B)}))\leq|B|\ln\nu\,.$ (4.2)
A bound on the difference of the two entropies is provided by the following variant of Fannes’ continuity bound [29, 11].
For the reader’s convenience, a proof is included here in Appendix A.
Under this relation, closeness in entropy requires proximity of the states in the trace norm to a degree which is affected by rank considerations. For concreteness sake let us consider the case where the following condition is met. (An example of that are the ground states of the quantum Ising model throughout its subcritical phase).
Definition 4.2.
A state $|\psi\rangle$ is said to exhibit exponentially fast conditional decoupling, at correlation length $\xi>0$, if for any regular $A\subset W$ and buffer $B_{l}=\{u\in W\backslash A\ |\ d(u,A)\leq l\}$ of width $l$:
${\delta}_{B_{l}}(A\,|\,C_{l})+\vartheta_{B_{l}}(A|C_{l})\leq C_{\xi}\ L(A)^{2\alpha}\ e^{-l/\xi}$ (4.4)
with $C_{l}=W\setminus(A\cup B_{l})$, $C_{\xi},\alpha\in(0,\infty)$ and $L(A)$ the length of $A$ from Definition 4.1.
In that case, the combination of (4.2) and (4.3) yields:
Theorem 4.3.
For a state $|\psi\rangle$, with exponentially fast conditional decoupling, the entanglement entropy corresponding to any regular $A\subset W$ with $L(A)\geq 2$ satisfies
$S(\varrho_{A}(\psi))\leq C_{\xi,d}\,|\partial A|\ln L(A)$ (4.5)
at some constat $C_{\xi,d}<\infty$ which is independent of $A$.
Proof.
Aiming for an approximation of relatively low rank and high fidelity, let us set the buffer’s width to $l=2(1+\alpha^{\prime})\xi\ln L(A)$ with $\alpha^{\prime}>\alpha$. For this choice $S(\varrho_{A}(\psi_{(B_{l})}))\ \leq\ |B_{l}|\log\nu\leq C_{d}\ l\ |\partial A|\ln\nu\,$, by (4.1) and thus the claimed bound holds for the approximating state.
Arranging the terms of (4.3) by order of magnitude, we get
$S(\varrho_{A}(\psi))-S(\varrho_{A}(\psi_{(B_{l})}))\ \leq\ \tfrac{1}{2}\Delta_{l}^{(1)}|A|\ln\nu+O(1),$ (4.6)
with $\Delta_{l}^{(1)}:=\left\|\varrho_{A}(\Psi)-\varrho_{A}(\Psi_{B})\right\|_{1}$ and the term $O(1)$ is the collection of terms $\tfrac{1}{2}\Delta_{l}^{(1)}\big{[}1+\ln\big{(}2/\Delta_{l}^{(1)}\big{)}\big{]}$. Under the assumption (4.4) the bound (3.6) yields
$\Delta_{l}^{(1)}\ |A|\leq C_{\xi}^{1/2}c_{d}\ L(A)^{\alpha-\alpha^{\prime}}|\partial A|$ (4.7)
which is negligible for $\alpha^{\prime}>\alpha$. Hence, under the stated condition the difference in entropies is insignificant. ∎
Reviewing the argument one may note that the above bound is suboptimal in that measuring the entropy difference in terms of $\|\varrho_{A}(\psi)-\varrho_{A}(\psi_{(B)})\|_{1}$ ignores the fact that the states’ disparity may be significant mainly in areas close to the boundary of $A$.
We next derive an improved upper bound in which $\ln|A|$ is replaced by $\ln|\partial A|$, derived through a multiscale analysis. The correction is particularly significant for one dimension, in which case $|\partial A|$ is of order $O(1)$ and hence the log term in (4.5) may be dropped.
4.2 An improved multiscale analysis
Inspecting (1.1) one may note that a state’s entropy $S(\rho_{A})$ may be strongly affected by the eigenvalues at the low end of $\varrho_{A}$’s spectrum. Taking a cue from [1, 16, 8], the following entanglement estimate is derived by decomposing the entropy $S(\varrho_{A})$ into a sum of contributions from different spectral regions, and employing buffers $B_{l}$ of increasing width $l\in\mathbb{N}$ for probing the contributions from the spectrum’s lower scales.
The analysis is conveniently quantified by the distribution-type function
@@ 4.2 An improved multiscale analysis
where the sum is over the eigenvalues of $\varrho_{A}$ labeled in decreasing order. Of particular relevance is the rate at which $\tau_{\varrho_{A}}(N)$ vanishes as $N$ increases toward the dimension of the relevant Hilbert space. The following main lemma estimates this quantity using the fidelity estimate (3.6).
@@
$\tau_{\varrho_{A}}(N)=1-\max\left\{\mathrm{tr}\varrho_{A}P_{N}\ |\ P_{N}\,\mbox{orthogonal projection of $\mathrm{rank}\ P_{N}\leq N$}\right\}$ (4.10)
in the case $N=\nu^{|B|}$ with $P_{N}$ is the spectral projection onto the span of the vectors $|\phi_{A}(\boldsymbol{\sigma}_{B})\rangle$ with $\boldsymbol{\sigma}_{B}\in\Omega_{B}$, cf. (2.10). Since $P_{N}\otimes\mathbbm{1}_{BC}|\psi_{(B)}\rangle=|\psi_{(B)}\rangle$ and $\mathbbm{1}\geq|\psi_{(B)}\rangle\langle\psi_{(B)}|$, we have
$\mathrm{tr}\varrho_{A}(\psi)P_{N}=\langle\psi|P_{N}\otimes\mathbbm{1}_{BC}|\psi\rangle\geq\left|\langle\psi|\psi_{(B)}\rangle\right|^{2}.$ (4.11)
The stated inequality then follows by estimating $1-\left|\langle\psi|\psi_{(B)}\rangle\right|^{2}\leq 2\left|1-\langle\psi|\psi_{(B)}\rangle\right|$ and the fidelity bound of Theorem 3.1. ∎
The following generalizes Definition 4.2.
@@
The distance function $l_{0}$ plays the role of the length scale at which the $\varphi$-decay sets in and allows to absorb $A$-dependent prefactors as in (4.4). These prefactors could be independent of $A$, in which case $l_{0}$ is constant.
In this setting, we obtain the following general entanglement bound.
@@
by the monotonicity of $F(x)=x\left(\ln(a/x)+1\right)$ for $0\leq x\leq 1\leq a$. Collecting the contributions to $S(\varrho_{A})=S_{0}+\sum_{k\geq 1}S_{k}$ and noticing that $l_{0}+k$ does not exceed $L(A)$, since the total number of eigenvalues of $\varrho_{A}\equiv\varrho_{A}(\psi)$ is $\nu^{|A|}$, yields the result. ∎
In order to elucidate this general statement, let us add some comments and formulate a simple consequence. In case $A$ is a cube, it is regular and $L(A)$ is proportional to the cube’s edge length. Moreover, power law decay of $\varphi$ with $\alpha>2$ is enough to ensure the convergence of the series corresponding to the sums in (4.14). One may therefore upper bound these sums independently of $A$. We thus arrive at the following, which is one of the main results of this work.
@@
For any state $|\psi\rangle$ which is conditionally $\varphi$-decoupled with power law decay with $\alpha>2$ beyond distance $l_{0}$, there is $C\in(0,\infty)$, independent of $l_{0}$, such that the state’s entanglement entropies satisfy
$S(\varrho_{A}(\psi))\leq C\ |\partial A|\ (1+l_{0}(A))\,$ (4.17)
for all regular $A\subset W$.
@@
The claim then follows from Theorem 4.6, with the additive constants absorbed into the multiplicative constant in the right side of (4.17). ∎
Several remarks are in order:
@@
    The bound (4.17) describes a strict area law in case the decoupling distance $l_{0}(A)$ is independent of $A$. This applies in particular to stoquastic states with $p$ any (even critical) classical Gibbs probability measure corresponding to a finite range interaction, cf. (1.8). Such states fall into the category of tensor-network states, for which the area-law of the entanglement entropy is well known (cf. [53, 55]).